OpenStudy (anonymous):

A blood test to detect prostate cancer in men gives a positive result 96% of the time if a person has prostate cancer, and it is 97% accurate for people who do not have the disease. What is the probability of getting a false positive result (that is, a person tests positive but does not actually have the ailment)? A.0.01 B.0.03 C.0.04 D.0.07 E.0.97

OpenStudy (anonymous):

@amistre64

OpenStudy (ivettef365):

if 97% is accurate then 3% could be the possible mistake

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

idk that sounds reasonable but its not an answer choice

OpenStudy (ivettef365):

yes it is, 3% equals to .03

OpenStudy (ivettef365):

to convert percent to decimal you just move the point twice to the left

OpenStudy (anonymous):

ohh, i never knew that..

OpenStudy (whpalmer4):

\[n\% = \frac{n}{100}\]

OpenStudy (ivettef365):

and to convert decimal to percent just move the point twice to the right example 0.97 = 97%

OpenStudy (anonymous):

what is n?

OpenStudy (ivettef365):

n is just a variable for any percent amount, for example n% = n/100 97% = 97/100

OpenStudy (anonymous):

idk are you sure its .03?

OpenStudy (ivettef365):

why are you not sure, what is it that you don't understand

OpenStudy (anonymous):

why isnt it .04?

OpenStudy (ivettef365):

because the question ask for the people that does NOT have the disease and get positive result, therefore you are working with the 97%

OpenStudy (anonymous):

alright ill go with .03

OpenStudy (ivettef365):

now if the question was what is the probability of a person having the disease and getting a not accurate result then you work with the 96%

OpenStudy (amistre64):

how many people do not have the ailment?

OpenStudy (anonymous):

what?

OpenStudy (amistre64):

97% accurate for people who do not have the disease. would indicate that its 3% false positive

OpenStudy (anonymous):

yea and then ivette said you move the decimal to the left twicce which gives me .03

OpenStudy (amistre64):

ivette was correct ... see, there are others :)

OpenStudy (anonymous):

yea i guesss..